OpenStudy (anonymous):

Alec, Hayden and Rose are comparing how much they get paid. Hayden gets £4 per hour more than Alec. The ratio of Rose's hourly rate to Alec's is 3:2 The total of their hourly rates of pay is £25. Which of the three earns the most? Show how you decide.

OpenStudy (anonymous):

So let income of Alec = A , Hayden = B, Rose = C Now Hayden gets 4 per hr more than Alec. Mathematically, we have - B = A + 4 Also rate of Rose to Alec = 3 : 2 Hence C : A = 3 : 2 Total hourly rate is 25. A+B+C = 25 A+A+4+C = 25 ( B = A + 4) 2A+C+4 = 25 => 2A+C = 21 Let C = 3x, A = 2x. (since C : A = 3:2) Using above values 2.2X + 3X = 21 7X = 21 => X = 3 Hence pay of A = 2x3 = 6, C = 3x3 = 9 B = 25-6-9 = 10 Hence, Alec gets 6 , Hayden gets 10, Rose gets 9. Hope it helps! :)

OpenStudy (anonymous):

Thanks.